OpenStudy (anonymous):

Solve the initial value problem: y' = x^3(1-y) y(0)=3

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{1}{1-y} dy= \int\limits_{}^{}x^{3} dx\]

OpenStudy (anonymous):

solve by seperation of variables now use a u=1-y substitution and you should be able to finish it

OpenStudy (anonymous):

\[u=1-y--------\int\limits \frac{1}{u} du= \ln(u) =\int\limits x^{3}=-\frac{x^{4}}{4}+C\]

OpenStudy (anonymous):

\[1-y = e^{- \frac{x^{4}}{4}+c}\]

myininaya (myininaya):

gj :)

OpenStudy (anonymous):

Thanks